# Quantum Lamb model

### v2

@@ I Introduction
Advances in the fabrication and characterization of simple mechanical systems in the nanoscopic or mesoscopic regime have driven experimental and theoretical investigations [1, 2] into some of the foundational principles of quantum mechanics. Prominent examples of such systems include vibrating beams and mirrored surfaces that interact with laser light through its radiation pressure (optomechanics) [3], mechanical resonators coupled to electronic devices (nanoelectromechanics) [4, 5], and interacting mechanical resonators (quantum acoustodynamics) [6, 7, 8]. In addition to providing a path to explore quantum science and the limits of precision measurement, such systems might be used to fashion new quantum sensors and devices for manipulating quantum information [9, 10].
@@
There have been other formulations of the damped quantum harmonic oscillator. Feshbach and Tikochinsky [12] introduced an auxiliary variable into the lagrangian of a harmonic oscillator to get the desired effective equation of motion for the damped oscillator. They then proceed by canonical quantization to obtain a quantum description of the damped harmonic oscillator. The auxiliary variable presumably functions as a single, effective environmental degree of freedom, but the connection to the microscopic physics is not made.
Caldeira and Leggett [13] separate the system into a sum of two subsystems (oscillator and bath) plus an interaction. Using a path integral description, the bath degrees of freedom can be integrated out to give a general quantum formulation of dissipative systems. Yurke [14] specifically considered a Lamb-type model that is a special case of the model considered here. (We will recover Yurke’s results by allowing the spring that couples the bead motion to the string to be suitably stiff.) Yurke considered a string with a point mass at one end. The point mass is also coupled to a spring with a fixed end. The mass-loaded string then has a time-dependent boundary condition. As a result, the normal modes are nonorthogonal. Yurke overcomes this by finding an appropriate weighting factor to use in redefining the inner product so that generalized orthogonality can be applied. He then quantized the model in the standard way.
@@
We show that the ground state of the quantum Lamb model is a nonclassical state–a multimode squeezed vacuum state– and we relate this ground state to the uncoupled states (transverse phonons of the string and vibrons of the bead) of the system. We then study the dynamics of the vibrational decay of the bead. We obtain an explicit expression for the decay rate, and we calculate the spectral distribution of single bogoliubon emission, a product of the decay. We show that for weak coupling strength $g$, the decay rate calculated agrees with both the classical result and the golden rule.
@@ II Hamiltonian
where $a_{n}^{\dagger}$ ($a_{n}$) creates (annihilates) a transverse acoustic phonon on the string, and $a_{0}^{\dagger}$ ($a_{0}$) creates (annihilates) a vibron on the bead. (We use index notation where (greek) $\alpha=0,1,2,\dots$, while (roman) $n=1,2,\dots,N$, and work with natural units where $\hbar=1$.)
@@
$\gamma_{n}=\frac{\kappa_{c}}{\sqrt{2\sigma m\omega_{0}\omega_{n}\ell}}\frac{k_{n}}{\sqrt{k_{n}^{2}+(\frac{\kappa_{c}}{\tau})^{2}}}\frac{1}{\sqrt{1-\frac{\kappa_{c}\ell/\tau}{(k_{n}\ell)^{2}+(\kappa_{c}\ell/\tau)^{2}}}}$ (2)
where $k_{n}$ is a solution of the transcendental equation $\tan k_{n}\ell=-\frac{\tau}{\kappa_{c}}k_{n}$.
@@ III Bogoliubov transformation
where the Bogoliubov frequencies $\{\Omega_{\alpha}\}$ satisfy the following nonlinear equation:
@@
By using the pole expansion form (Mittag-Leffler) for cotangent, it is straightforward to show that in the thermodynamic limit ($\ell,N\to\infty$ and $\frac{N}{\ell}\to\frac{\omega_{d}}{\pi c}$), Eq. 10 approaches Yurke’s transcendental equation [14] for the special coupling case of $\omega_{c}=\omega_{c}^{*}\equiv\sqrt{\frac{4}{\pi}\nu\omega_{d}}$
@@
The frequency $\omega_{d}$ is recognized as the Debye frequency of the string (the high-frequency cutoff for the string) and $\nu=\frac{\tau}{2mc}$, the classical damping rate for the bead oscillation amplitude.
@@
We define the coupling strength $g\equiv\frac{4}{\omega_{0}}\sum_{n}{\frac{\gamma_{n}^{2}}{\omega_{n}}}$ and conclude that there is a critical coupling strength $g_{c}=1$ above which the model is ill-defined. Using the form for $\gamma_{n}$ in Eq. 2, $g$ can be approximated by $g\approx\frac{4\nu\omega_{d}}{\pi\omega_{c}^{2}}$ for suitably stiff $\kappa_{c}$ ($\omega_{c}^{2}\gg 2\nu\omega_{d}$). Thus, $g\ll\frac{2}{\pi}$ in this regime, and the model is stable. The model is also stable for the case of $\omega_{c}=\omega_{c}^{*}$, since the behavior of Eq. 11 near $\Omega_{\alpha}\sim 0$ reveals that a zero frequency solution does not exist.
For low $\omega_{c}$ ($\omega_{c}^{2}\ll\frac{2\pi c\nu}{\ell}$), $g\ll\frac{2}{\pi}\frac{\omega_{c}^{2}}{\omega_{b}^{2}+\omega_{c}^{2}}$. Hence, the model is stable for suitably low $\kappa_{c}$. We found numerically that the coupling strength $g$ can take values close to 1 (see Fig. 2). Whether $g$ can exceed 1 in the physical parameter space remains an open question.
Figure 2: Coupling strength $g$ versus $\nu/\omega_{0}$. Maximum value $g=0.995$ for parameter values $\nu/\omega_{0}=0.56$, $\omega_{c}/\omega_{b}=22.2$ and $\omega_{d}/\omega_{c}=300$.
@@ IV Multimode squeezed vacuum
Hence, the ground state of the model is a multimode squeezed vacuum state [19, 16].
The average number of uncoupled excitations (phonons and vibrons) in the mode $\alpha$ contained in the coupled ground state |$|\{0\}\rangle$ can be expressed in terms of Bogoliubov coefficients, with
@@ VI Summary
Our solution explicitly calculates the coefficients of the multimode Bogoliubov transformation that diagonalizes the Hamiltonian, and we use these coefficients to describe the properties of the system. We found that the true ground state of the system is a multimode squeezed vacuum state, and we obtained the bogoliubon spectrum. We then obtained an explicit expression for the vibrational decay rate of the bead and found that it recovered the classical damping rate in the light damping regime.
We examined the acoustic radiation spectrum emitted by the vibrating particle. We obtained an expression for the probability of single bogoliubon emission in terms of the the Bogoliubov coefficients, and we observed that the spectral emission has a nearly symmetric lineshape about a slightly red-shifted peak frequency.

### v3

@@ I Introduction
Advances in the fabrication and characterization of simple mechanical systems in the nanoscopic and mesoscopic regime have facilitated experimental and theoretical investigations [1, 2] into some of the foundational principles of quantum mechanics. Prominent examples of such systems include vibrating beams and mirrored surfaces that interact with laser light through its radiation pressure (optomechanics) [3], mechanical resonators coupled to electronic devices (nanoelectromechanics) [4, 5], and interacting mechanical resonators (quantum acoustodynamics) [6, 7, 8]. In addition to providing a path to explore quantum science and the limits of precision measurement, such systems might be used to fashion new quantum sensors and devices for manipulating quantum information [9, 10].
@@
There have been other formulations of the damped quantum harmonic oscillator. Feshbach and Tikochinsky [12] introduced an auxiliary variable into the lagrangian of a harmonic oscillator to get the desired effective equation of motion for the damped oscillator. They then proceeded by canonical quantization to obtain a quantum description of the damped harmonic oscillator. The auxiliary variable presumably functions as a single, effective environmental degree of freedom, but the connection to the microscopic physics is not made.
Caldeira and Leggett [13] separate the system into a sum of two subsystems (oscillator and bath) plus an interaction. Using a path integral description, the bath degrees of freedom can be integrated out to give a general quantum formulation of dissipative systems. Yurke [14] specifically considered a Lamb-type model that is a special case of the model considered here. (We will recover Yurke’s results by allowing the spring that couples the bead motion to the string to be suitably stiff.) Yurke considered a string with a point mass at one end. The point mass is also coupled to a spring with a fixed end. The mass-loaded string then has a time-dependent boundary condition. As a result, the normal modes are nonorthogonal. Yurke overcame this by finding an appropriate weighting factor to use in redefining the inner product so that generalized orthogonality can be applied. He then quantized the model in the standard way.
@@
We show that the ground state of the quantum Lamb model is a nonclassical state–a multimode squeezed vacuum state– and we relate this ground state to the uncoupled states (transverse phonons of the string and vibrons of the bead) of the system. Squeezed states can serve as a quantum resource for precision sensing applications; for example, gravitational wave detection relies on squeezing to perform displacement measurements where uncertainty in momentum is sacrificed in favor of reduced uncertainty in position. Caldeira and Leggett [13] found in their studies of the damped quantum oscillator, the uncertainty in position for the ground state is reduced with increasing damping, strongly so in the overdamped regime. This result is consistent with a squeezed ground state.
We then study the dynamics of the vibrational decay of the bead. Dissipation emerges in the thermodynamic limit where the number of string modes $N$ becomes infinitely large ($N\to\infty$). In this limit, vibrational energy of bead can be radiated away. We then obtain an explicit expression for the decay rate, and we calculate the spectral distribution of single bogoliubon emission, a product of the bead decay. We show that for weak coupling strength $g$, the decay rate calculated agrees with both the classical result and the golden rule.
@@ II Hamiltonian
where $a_{n}^{\dagger}$ ($a_{n}$) creates (annihilates) a transverse acoustic phonon on the string, and $a_{0}^{\dagger}$ ($a_{0}$) creates (annihilates) a vibron on the bead. (We use index notation where (greek) $\alpha=0,1,2,\dots$, while (roman) $n=1,2,\dots,N$, and work with natural units where $\hbar=1$.) $N$ is the number of vibrational modes for the string. We are ultimately interested in the limit $N\to\infty$ in order to obtain a description of the damped bead oscillator.
@@
$\gamma_{n}=\omega_{s}\sqrt{\frac{\nu}{\omega_{0}}}\sqrt{\frac{k_{n}\ell}{(k_{n}\ell)^{2}+(k_{s}\ell)^{2}}}\frac{1}{\sqrt{1-\frac{k_{s}\ell}{(k_{n}\ell)^{2}+(k_{s}\ell)^{2}}}}$ (2)
where $\omega_{s}=ck_{s}=\frac{c\kappa_{c}}{\tau}$ and $\nu=\frac{\tau}{2mc}$. The wavenumber $k_{n}$ is a solution of the transcendental equation $\tan k_{n}\ell=-\frac{\tau}{\kappa_{c}}k_{n}$.
@@ III Bogoliubov transformation
where the Bogoliubov frequencies $\{\Omega_{\alpha}\}$ satisfy the following summation equation:
@@
By using the pole expansion form (Mittag-Leffler) for cotangent, it is straightforward to show that in the thermodynamic limit ($\ell,N\to\infty$ and $\frac{N}{\ell}\to\frac{\omega_{d}}{\pi c}$), Eq. 10 gives Yurke’s transcendental equation [14] for the special coupling case of $\omega_{c}=\omega_{c}^{*}\equiv\sqrt{\frac{4}{\pi}\nu\omega_{d}}$
@@
The frequency $\omega_{d}$ is recognized as the Debye frequency of the string (the high-frequency cutoff for the string).
@@
We define the coupling strength $g\equiv\frac{4}{\omega_{0}}\sum_{n}{\frac{\gamma_{n}^{2}}{\omega_{n}}}$ and from Eq. 12 conclude that there is a critical coupling strength $g_{c}=1$ above which the model is ill-defined. Using the form for $\gamma_{n}$ in Eq. 2, we obtain an expression for $g$ in the thermodynamic limit
$g=\frac{2}{\pi}\bigg{(}\frac{\omega_{c}}{\omega_{0}}\bigg{)}^{2}\tan^{-1}\frac{\pi\tau}{\kappa_{c}d}$ (13)
where $d=\ell/N$, the interatomic distance between atoms in the string. With increasing string tension $\tau$, $g$ asymptotically approaches $g=\big{(}\frac{\omega_{c}}{\omega_{0}}\big{)}^{2}$, a quantity bounded by 1 (see Fig. 2). Thus, the model is stable over the range of physical parameters. (As a practical matter, before the tension $\tau$ becomes comparable to interatomic forces in the string, it is likely that the string would break.)
Figure 2: Coupling strength $g$ versus $\frac{\tau}{\kappa_{c}d}$ for $N,\ell\to\infty$.
@@ IV Multimode squeezed vacuum
Hence, the ground state of the model is always a multimode squeezed vacuum state [19, 16] with squeeze parameter $\xi$ determined by the Bogoliubov coefficients.
We note that the position uncertainty of the bead in the ground state can be readily evaluated with the Bogoliubov coefficients. Evaluating the bead variance gives
$\langle\{0\}|u_{0}^{2}|\{0\}\rangle$ $=$ $\frac{\hbar}{2m\omega_{0}}\sum_{\alpha}(M_{\alpha 0}-N_{\alpha 0})^{2}$ (16)
$=$ $\frac{\hbar}{2m}\sum_{\alpha}\frac{1}{\Omega_{\alpha}(1+4\omega_{0}\sum_{q}\frac{\gamma_{q}^{2}\omega_{q}}{(\Omega_{\alpha}^{2}-\omega_{q}^{2})^{2}})}$ (17)
The sum can be evaluated analytically using complex contour integration [15] to reveal that the position uncertainty is reduced with increasing damping $\nu$, a result first obtained by Caldeira and Leggett [13] using the fluctuation-dissipation theorem.
To better understand the ground state, the average number of uncoupled excitations (phonons and vibrons) in the mode $\alpha$ contained in the coupled ground state |$|\{0\}\rangle$ can also be expressed in terms of Bogoliubov coefficients, with
@@ VI Summary
Our solution explicitly calculates the coefficients of the multimode Bogoliubov transformation that diagonalize the Hamiltonian, and we use these coefficients to describe the properties of the system. We found that the true ground state of the system is always a multimode squeezed vacuum state where the displacement uncertainty is reduced with increasing damping rate $\nu$. We then obtained an explicit expression for the vibrational decay rate of the bead and found that it recovered the classical damping rate in the light damping regime.
We examined the acoustic radiation spectrum emitted by the vibrating particle. We obtained an expression for the probability of single bogoliubon emission in terms of the Bogoliubov coefficients, and we observed that the spectral emission has a nearly symmetric lineshape about a slightly red-shifted peak frequency.